OpenStudy (anonymous):

what is the r valueof 40,20,10,5

OpenStudy (anonymous):

what do you have to do to get from 40 to 20?

OpenStudy (zzr0ck3r):

what is 20/40=10/20=5/10?

OpenStudy (anonymous):

divide by 2

OpenStudy (anonymous):

yes divide by 2...which is the same thing as multiplying by what?

OpenStudy (anonymous):

2 and thanks

OpenStudy (anonymous):

no its the same thing as multiplying by \({1\over2}\)

OpenStudy (anonymous):

so your r-value=\(\large {1\over2}\)

OpenStudy (anonymous):

oh ok I get it thx

OpenStudy (zzr0ck3r):

take the second number and divide it by the first, take the third number divide it by the second, take the fourth number and divide by the third, they should all be the same number and that number is your answer

OpenStudy (anonymous):

np

OpenStudy (anonymous):

wow that makes it really easy thx

OpenStudy (anonymous):

is 3,4,16:9,64;27 a geometric sequence?

OpenStudy (zzr0ck3r):

no

OpenStudy (anonymous):

is 1,5,-25,125

OpenStudy (zzr0ck3r):

no but -1,5,-25,125 is...I think you get it

OpenStudy (zzr0ck3r):

ratio must stay constant

OpenStudy (anonymous):

ok yea I get it

OpenStudy (zzr0ck3r):

you could make sure by dividing the first term by the second, and make that is constant. But that does not give you the ratio you want, you want 2nd term divided by the first.

OpenStudy (anonymous):

ok